# Spectral Deconvolution without the Deconvolution: Extracting Temperature from X-ray Thomson Scattering Spectra without the Source-and-Instrument Function

### v1

@@ I Introduction
Figure 1: (a) Ray traced IFs on the spectrometer used in this work, measured with different single photon energies $E_{0}$ on the crystal: 8.1 keV (black), 8.3 keV (red), and 8.5 keV (blue). (b) Ray traced IFs at 8.5 keV with the crystal at different positions $\Delta l$ along the dispersion axis. In both plots, the sudden drops in the intensity correspond to the physical edges of a crystal in space being reached, so only part of the crystal reflects photons to the detector.
@@
X-ray Thomson scattering (XRTS) has emerged as a workhorse spectroscopy technique in the diagnosis of HED systems [26, 27, 28, 7, 4, 29, 2, 30, 31, 32, 33, 34, 35, 36]. By scattering photons of energy $E_{i}$ off the electrons in a system, the electronic dynamic structure factor (DSF) $S(\bm{q},E_{i}-E_{s})$ can be probed along a given scattering vector $\bm{q}$ and spectrally resolved in the scattered photon energy $E_{s}$. The scattering spectrum emerging from the target $I_{s}(\bm{q},E_{s})$ is the convolution of the DSF with the incident beam profile $B(E_{i})$,
@@
where $E$ is the photon energy as inferred by the spectrometer calibration, and $\bm{\Gamma}(E_{s})$ are the parameters that describe the broadening by the crystal which, for a given crystal, depends on the energy of the scattered photon [38, 39]. The IF also depends strongly on where the crystal is positioned in space [37, 40], as shown in Fig. 1. Evidently, the crystal IF does not broaden the spectrum as a convolution. Nevertheless, in modelling, the IF is often applied as a simple convolution to the spectrum, so the final measured spectrum is just treated as the DSF convolved with the combined source-and-instrument function (SIF) $R(E)$ of the setup:
@@
where the approximation $E_{s}/E_{i}\approx 1$ across the spectral range has also been made. Within a sufficiently limited spectral window, which depends on the widths of features being measured, this approximation produces spectra that look very similar to the full kernel equation from Eq. (2[37].
@@
One of the main practical challenges in using the ITCF method is the accurate characterization of the SIF [51, 37]. Much of the difficulty is in characterising the IF of the mosaic crystals that are widely employed in XRTS experiments. Recent works have shown their IFs can extend over the whole detected spectral range with features that range over several orders of magnitude in intensity [37, 40], shown in Fig. 1. Due to the strong geometry dependence of the IF – shown in particular by Fig. 1 (b) – it is necessary to measure the SIF in every experiment. One common approach for measuring the SIF directly is to measure the shape of the quasi-elastic scattering since, except for ultrahigh resolution setups [52, 53, 54], it just reproduces the SIF. However, eventually the quasi-elastic scattering will mix with inelastic scattering, meaning full characterization of the SIF is not possible. In the case of x-ray backlighters [55, 4, 56, 2, 43], there is also a strong dependency of the emission line intensities on the laser heating conditions for each shot. But, due to restrictions in the experimental setups, the actual emission spectrum for a given shot is rarely measured. Therefore, models or approximations of the SIF are used in place of the actual SIF, which reintroduces a model-dependency in the underlying model-free analysis.
@@
If the system is in thermal equilibrium, the two ITCFs are symmetric around the same minimum point, and therefore the ratio of the ITCFs will also be symmetric around this point, which may be a minimum or a maximum depending on the relative size and curvature of each ITCF (for the purposes of this work, we consider the minima of the indicated ratios). This means the temperature can be extracted directly from this ratio. With three or more scattering angles, comparisons of the ratios allows a consistency check: if the temperatures inferred are different between the ratios, or if the ratios themselves are not symmetric around their minimum or maximum, these would be key indicators of the system being out of equilibrium [48]. Crucially, by bypassing the explicit deconvolution of the SIF, this ratio method is truly model-free since no model of the DSF nor the SIF is required.
@@ II.1 Dynamic Structure Factors
Dynamic structure factors were calculated using the multi-component scattering simulation (MCSS) code [57, 58], which uses the Chihara decomposition [59, 60] to calculate the DSF. We simulated carbon with mass density 3.5 g/cc and a single C4+ ion in the charge state distribution. The ions and electrons are in equilibrium with one another at temperatures of 15, 20, 50, and 100 eV. The free-free feature was calculated using the random phase approximation (RPA) [61, 21], while the bound-free and free-bound features were calculated with the impulse approximation (IA) [62, 63]. At the resolution of the spectrometers considered here, the details of the ion feature cannot be resolved, so the elastic feature is treated as a delta-function in frequency space multiplied by the Rayleigh weight. The Rayleigh weight then just determines the portion of the photons which elastically scatter. The screening cloud within the Rayleigh weight is calculated using finite-wavelength screening and a Coulomb potential with the effective charge of $Z=4$ for the electron-ion potential. The ionic form factor uses the simple hydrogenic orbital model for its description. The static ion structure is calculated by a hyper-netted chain solver using a Debye-Hückel potential featuring a general screening length for the ion-ion interaction. The DSFs were calculated for a photon energy of 8.5 keV, and at scattering angles of 10, 20, and 150. We consider the spectrum incident on the spectrometer come from a single scattering vector, which is a reasonable approximation given the large incident photon energy compared to the size of the spectral window used in the deconvolution ($\le 250$ eV) [50].
@@ II.2 Ray Tracing
Based on fits to the quasi-elastic feature of polypropylene, the mosaic distribution of the crystallites was found to be well-described by a wrapped-Lorentzian-like distribution, with one of the crystals having a mosaic spread of 0.063 [40]. Previous measurements of the mosaic spread in HAPG suggest there is some dependence of the mosaic spread on the photon energy below 11 keV [39], however it is very weak and so here it is assumed to be constant for all photon energies.
@@
Since the IF dominates the complexity of the full SIF, we treat the probing beam as a Gaussian centred at 8.5 keV with a FWHM of 1 eV, similar to having the beam pass through a monochromator. The initial distribution of photons that are ray traced is given by Eq. (1). The broadening of the spectra is dominated by the crystal, which substantially alters the shape and deconvolution behaviour in the ratio method. Still, it should be noted that monochromators see frequent use at XFELs to improve the spectral resolution of the measured DSF [70], so our choice of source function is representative of realistic spectra collected at XFELs.
@@
Examples of the ray traced spectra are shown in Fig. 3. Around 8.55 keV and 8.68 keV, there are sudden drops in the intensity of the elastic signal corresponding to the physical edges of the crystal in space. The shape of the elastic feature has a noticeable effect on the shape of the spectra for the low scattering angles and the low temperatures, where features are narrowest. There is also a strong degree of asymmetry and a very extended high energy tail in the SIF due to the crystal having a Lorentzian mosaic distribution function [37]. On the downshifted side, there is also an extended tail due to the IRC. At high scattering angles for $T=50$ and 100 eV, the effect of the shape of the SIF is hidden by the very broad inelastic features. Taken together, the extended nature of this SIF and its complexity can be challenging to measure, and deconvolving it from XRTS spectra effectively could be challenging.
Finally, the Laplace transform expressed in Eq. (4) is centred symmetrically around $\omega=0$, which means a central photon energy needs to be defined. The choice is important since, if an error $\Delta E$ is made, an additional factor of $\exp(-\tau\Delta E)$ appears in front of the ITCF which will skew the position of its minimum. Even though the central beam energy is 8.5 keV, this is not the appropriate choice due to the asymmetric nature of the SIF. Instead, the SIF intensity-weighted average of the spectrometer energy is used to define the central energy of the scattering spectrum. If the SIF were symmetric, then this would align with the central photon energy of the beam.
@@ III.1 Checking the Convolution Theorem Approximation
As mentioned previously, the ITCF method relies on the approximation that the SIF is applied as a convolution [44, 50], however this is not strictly true in the case of the crystal IF [37] and this approximation has yet to be confirmed. Given the ratio method proposed here also relies on this approximation, we first test whether the temperature can be extracted using the convolution theorem.
@@
In all cases, once the SIF is deconvolved from the spectrum, a region of convergence in the inferred temperature with respect to $x$ is observed. Oscillations in the convergence may be taken as an uncertainty in the inferred temperature. For all cases considered here, the temperature converges to the expected value within $\lesssim 10$%. The inferred temperature for 150 at 15 eV is on the -10% limit, and this underestimation is likely due to the very weak upshifted inelastic signal for this spectrum (see Fig. 3), which would make it difficult to correctly probe detailed balance. However, this still only represents an absolute error of 1.5 eV, which is comparable to the absolute error observed at the other temperatures.
@@
In general, there is a small but seemingly systematic underestimation of the temperatures extracted here. This is likely due to the solid angle coverage of the crystal, which is larger on the parts of the crystal closer to the source, which corresponds to lower photon energies. In other words, the spectrometer has a higher collection efficiency at lower photon energies. This means that it appears as if there are more downshifted photons than there are, which skews the apparent detailed balance. At the same time, the $(E_{s}/E_{i})^{2}$ factor in Eq. (1) term mostly counteracts this effect by increasing the number of upshifted photons than detailed balance otherwise implies. While these effects will not be explicitly handled here since they mostly cancel, future improvements in analysis could account for them.
@@
Otherwise, the stability of the convergences and accuracy of the inferred temperature is very promising for the ITCF method in being able to accurately extract system conditions, particularly in comparison to forward modelling where error bars of up to 50% have been previously reported  [24, 7, 4]. This good performance is despite the SIF not strictly applying as a deconvolution to the DSF due to the spectrometer response function, Eq. (2), and the $(E_{s}/E_{i})^{2}$ factor in Eq. (1). To within at least $\pm 250$ eV around the elastic feature, the convolution approximation appears to hold very well.
@@
For the $T=15$ eV and 20 eV cases, sharp drops of the inferred temperature to $T_{\rm exact}/2$ are eventually seen as $x$ increases – such behaviour will also be seen later when using the ratios of ITCFs. These correspond to the ITCF losing its minimum point as the upshifted signal becomes very weak and few, if any, photons are measured. This causes the $\tau>\beta/2$ side of the ITCF to collapse, at which point the ITCF minimum just becomes the upper limit of the considered $\tau$-range, here restricted to $\tau_{\rm max}\ =\beta$. Such drops are not seen for $T=50$ eV and 100 eV since there is strong upshifted signal up to the edge of the detector. These drops in the inferred temperature indicate the limits of being able to perform the deconvolution and extract temperature due to lack of upshifted signal. It also demonstrates that the lower temperature limit that the ITCF method can measure is ultimately determined by the photon statistics: enough signal needs to be measured sufficiently far from the elastic feature in both the upshifted and downshifted sides of the spectrum so that convergence in $x$ can be observed. The more extended the SIF is, the further away this point will be due to the need to deconvolve a broader function. And at the same time, the detailed balance relationship means the number of upshifted photons is exponentially reduced as the temperature decreases.
@@
Likewise, the ratio of inelastic to elastic scattering matters since it is the inelastic signal which needs to be measured in order to probe detailed balance. To emphasise this point, Fig. 5 shows the convergence of the inferred temperature for $T=15$ eV when the Rayleigh weight has been reduced by a factor of 4, which increases the portion of photons scattered inelastically. In this case, it takes a larger integration range $x$ before the ITCF loses its minimum, and the agreement between the temperatures at the different angles is improved. Convergences for a given number of input photons therefore have a strong dependence on the Rayleigh weight, and the models in MCSS are known to often overestimate the elastic scattering contribution [57, 58]. In the remainder of the article, the standard DSF model is used for the simulations, but will be compared to results for when the Rayleigh weight is reduced as well to understand the role of the elastic feature in extracting the temperature.
@@ III.2 Temperature from ITCF Ratios
For the $T=20$ eV, 50 eV, and 100 eV cases, there are three ratios which show convergence, all with the higher scattering angle in the numerator of the ITCF ratio. The $T=50$ eV and 100 eV cases in particular produce very accurate temperatures over a very large range of $x$ due to their large upshifted signal. We also note these two high temperatures show more accurate temperature inferences using the ratio method in Fig. 6 than using the standard ITCF method in Fig. 4. The other ratios do not show meaningful, if any, convergence – however, these can simply be discarded since their inverses do show convergence (or equivalently, one could look for their maximum points). For $T=20$ eV, at $x>160$ eV, there are drops in the inferred temperature to $T_{\rm exact}/2$ which, as with ITCFs for Fig. 4, indicates the ITCF ratio has lost its well-defined minimum due to the $\tau>\beta/2$ side collapsing because of the lack of measured upshifted photons.
@@
For $T=15$ eV, only one ITCF ratio, 201020/10, shows a region of convergence in its minimum that could be used to extract temperature; the remainder do not yield useful information. This is because of the very low number of upshifted photons measured. Indeed, the situation is worse than in Fig. 4 since we now require upshifted signal in two ITCFs in order to measure a temperature. If the Rayleigh weight is reduced by a factor of 4, the ratio method is more successful at extracting the temperature, a shown in Fig. 7. In this case, there are three ratios – 2015020/150, 15010150/10, and 201020/10 – which have a region of convergence in their minima that can infer the temperature. When $x>140$ eV, these minima of these ratios begin to diverge or collapse to $T_{\rm exact}/2$. The 15 eV results again highlight that the ability to measure low temperatures is strongly determined by the number of measured upshifted photons, which from detailed balance becomes exponentially more difficult as the temperature decreases. For the HAPG crystal used here, the problem is made worse by the extended crystal IF, meaning upshifted photons need to be measured relatively far from the elastic peak in order to observe convergence. Low-mosaicity HOPG crystals may therefore be more useful for low temperature measurements since the Gaussian distribution of crystallites reduces the extent of the crystal broadening [39].
@@
Figs. 6 and 7 also highlight the strength of collecting XRTS spectra at three or more scattering angles: the analysis of the convergence is made much simpler by being able to cross-check the extracted temperatures between three or more ratios. At the same time, for the T=20T=20–100 eV cases, the high consistency in the inferred temperatures suggests that discrepancies of even a few eV in the inferred temperatures from the ratios of three or more scattering angles would be a strong indication of non-equilibrium effects in the system [48, 49]. However, as indicated by the $T=15$ eV results, the ability to do so is closely tied to the number of measured photons, and this caveat is explored in more detail in the following Section III.3.

### v2

@@ I Introduction
Figure 1: (a) Ray traced IFs on the spectrometer used in this work, which are representative of typical IFs for HAPG von Hámos spectrometers. The IFs are measured with different single photon energies $E_{0}$ on the crystal: 8.1 keV (black), 8.3 keV (red), and 8.5 keV (blue). (b) Ray traced IFs at 8.5 keV with the crystal at different positions $\Delta l$ along the dispersion axis. In both plots, the sudden drops in the intensity correspond to the physical edges of a crystal in space being reached, so only part of the crystal reflects photons to the detector (see Ref. [32] for further details).
X-ray Thomson scattering (XRTS) has emerged as a workhorse spectroscopy technique in the diagnosis of HED systems [44, 39, 57, 54, 16, 70, 28, 29, 15, 69, 19, 30, 22, 3]. By scattering photons of energy $E_{i}$ off the electrons in a system, the electronic dynamic structure factor (DSF) $S(\bm{q},E_{i}-E_{s})$ can be probed along a given scattering vector $\bm{q}$ and spectrally resolved in the scattered photon energy $E_{s}$. The scattering spectrum emerging from the target $I_{s}(\bm{q},E_{s})$ is the DSF broadened by the incident beam profile $B(E_{i})$,
@@
where $E$ is the photon energy as inferred by the spectrometer calibration, and $\bm{\Gamma}(E_{s})$ are the parameters that describe the broadening by the crystal. Typically, these parameters depend on the energy of the scattered photon [79, 36]. The IF also depends strongly on where the crystal is positioned in space [32, 35], as shown in Fig. 1. Evidently, the crystal IF does not broaden the spectrum as a convolution. Nevertheless, in practice, the IF broadening is often treated as a simple convolution to the spectrum. The final measured spectrum is then just treated as the DSF convolved with the combined source-and-instrument function (SIF) $R(E)$ of the setup:
@@
where the approximation $E_{s}/E_{i}\approx 1$ across the spectral range has also been made.
@@
One of the main practical challenges in using the ITCF method is the accurate characterization of the SIF [61, 32]. Much of the difficulty is in characterising the IF of the mosaic crystals that are widely employed in XRTS experiments. Recent works have shown their IFs can extend over the whole detected spectral range with features that range over several orders of magnitude in intensity [32, 35]; see Fig. 1. Due to the strong geometry dependence of the IF – shown in particular by Fig. 1 (b) – it is necessary to measure the SIF in every experiment.
One common approach for measuring the SIF directly is to measure the shape of the quasi-elastic scattering since, except for ultrahigh resolution setups [64, 78, 14], it just reproduces the SIF. However, eventually the quasi-elastic scattering will mix with inelastic scattering, meaning full characterization of the SIF is not possible. In the case of x-ray backlighters [66, 16, 17, 28, 56], there is also a strong dependency of the emission line intensities on the laser heating conditions for each shot. But, due to restrictions in the experimental setups, the actual emission spectrum for a given shot is rarely measured. Therefore, models or approximations of the SIF are used in place of the actual SIF, which reintroduces a model-dependency in the underlying model-free analysis.
@@
If the system is in thermal equilibrium, the two ITCFs are symmetric around the same minimum point, and therefore the ratio of the ITCFs will also be symmetric around this point. For the ratios, this may be a minimum or a maximum depending on the relative size and curvature of each ITCF, so for the purposes of this work we consider the minima of the indicated ratios. This means the temperature can be extracted directly from this ratio. With three or more scattering angles, comparisons of the ratios allows a consistency check: if the temperatures inferred are different between the ratios, or if the ratios themselves are not symmetric around their minimum or maximum, these would be key indicators of the system being out of equilibrium [77]. Crucially, by bypassing the explicit deconvolution of the SIF, this ratio method is truly model-free since no model of the DSF nor the SIF is required.
@@ II.1 Dynamic Structure Factors
Dynamic structure factors were calculated using the multi-component scattering simulation (MCSS) code [6, 5], which uses the Chihara decomposition [7, 8] to calculate the DSF. We simulated carbon with mass density 3.5 g cm-3 and a single C4+ ion. The ions and electrons are in equilibrium with one another at temperatures of 15, 20, 50, and 100 eV. The free-free feature was calculated using the random phase approximation (RPA) [37, 23], while the bound-free and free-bound features were calculated with the impulse approximation (IA) [27, 63]. At the resolution of the spectrometers considered here, the details of the ion feature cannot be resolved, so the elastic feature is treated as a delta-function in frequency space multiplied by the Rayleigh weight. The Rayleigh weight then just determines the portion of the photons which elastically scatter. The screening cloud within the Rayleigh weight is calculated using finite-wavelength screening and a Coulomb potential with the effective charge of $Z=4$ for the electron-ion potential. The ionic form factor uses the simple hydrogenic orbital model for its description. The static ion structure is calculated by a hyper-netted chain solver using a Debye-Hückel potential featuring a general screening length for the ion-ion interaction. The DSFs were calculated for a photon energy of 8.5 keV, and at scattering angles of 10, 20, and 150. We consider the spectrum incident on the spectrometer come from a single scattering vector, which is a reasonable approximation given the large incident photon energy compared to the size of the spectral window used in the deconvolution ($\le 250$ eV) [21].
@@ II.2 Ray Tracing
Based on fits to the quasi-elastic feature of polypropylene, the mosaic distribution of the crystallites was found to be well-described by a wrapped-Lorentzian-like distribution, with one of the crystals having a mosaic spread of 0.063 [35]. Previous measurements of the mosaic spread in HAPG suggest there is some dependence of the mosaic spread on the photon energy below 11 keV [36], however the dependence is very weak, and so here it is assumed to be constant for all photon energies.
@@
Since the IF dominates the complexity of the full SIF, we treat the probing beam as a Gaussian centred at 8.5 keV with a FWHM of 1 eV. This is similar to having the beam pass through a monochromator. The initial distribution of photons that are ray traced is given by Eq. (1). The broadening of the spectra is dominated by the crystal, which substantially alters the shape and deconvolution behaviour in the ratio method. Still, monochromators see frequent use at XFELs to improve the spectral resolution of the measured DSF [75, 34, 33], so our choice of source function is representative of realistic spectra collected at XFELs.
@@
The simulation setup described above is most relevant for XFEL experiments. To demonstrate the applicability of the method to a broader range of experiments, in Section III.6 we consider a NIF-style experiment based on Ref. [16]. Further details on those simulations may be found in that section.
The detector images from the ray tracing simulations are output in terms of number of photons per pixel, so the resulting spectra consider number of detected photons per energy bin. While HEART is able to model filter and the quantum efficiency effects, usually these are either accounted for when processing a spectrum, or they are sufficiently uniform in the spectral range to be ignored. Additionally, for the detectors at XFELs, their response is typically calibrated ahead of experiments with known sources, allowing for a conversion of the analogue-to-digital units into the total photon energy deposited in a pixel. This can then be converted to photon counts via the spectrometer dispersion. Provided the spectrum units are proportional to photon counts per energy bin, the specific units are otherwise arbitrary since this just gives a constant in front of the ITCF. The ITCF could be normalised to its true units [18], but this is unnecessary for the anaylsis performed here.
Examples of the ray traced spectra are shown in Fig. 3. Around 8.55 keV and 8.68 keV, there are sudden drops in the intensity of the elastic signal corresponding to the physical edges of the crystal in space (see Ref. [32] for further details). The shape of the elastic feature has a noticeable effect on the shape of the spectra for the low scattering angles and the low temperatures, where features are narrowest. At high scattering angles for $T=50$ and 100 eV, the effect of the shape of the SIF is hidden by the very broad inelastic features. Taken together, the extended nature of this SIF and its complexity can be challenging to measure, and deconvolving it from XRTS spectra effectively could be challenging.
The Laplace transform expressed in Eq. (4) is centred symmetrically around $\omega=0$, which means a central photon energy needs to be defined. The choice is important since if an error $\Delta E$ is made, an additional factor of $\exp(-\tau\Delta E)$ appears in front of the ITCF which will skew the position of its minimum. Even though the central beam energy is 8.5 keV, this is not the appropriate choice due to the asymmetric nature of the SIF. Instead, the SIF intensity-weighted average of the spectrometer energy is used to define the central energy of the scattering spectrum.
Finally, the convergence of the inferred temperatures from the figures is identified visually, given the large variety in shape and form of the curves.
@@ III.1 Checking the Convolution Theorem Approximation
As mentioned previously, the ITCF method relies on the approximation that the SIF is applied as a convolution [20, 21]. However, this is not strictly true in the case of the crystal IF, and this approximation has yet to be confirmed. Given the ratio method proposed here also relies on this approximation, we first test whether the temperature can be extracted using the convolution theorem.
@@
In all cases, once the SIF is deconvolved from the spectrum, a region of convergence in the inferred temperature with respect to $x$ is observed. Oscillations in the convergence may be taken as an uncertainty in the inferred temperature. For all cases considered here, the temperature converges to the expected value within $\lesssim 10$%. The inferred temperature for 150 at 15 eV is on the -10% limit, and this underestimation is likely due to the very weak upshifted inelastic signal for this spectrum (see Fig. 3), so it is difficult to probe detailed balance. This still only represents an absolute error of 1.5 eV, which is comparable to the absolute error observed at the other temperatures.
@@
In general, there is a small but seemingly systematic underestimation of the temperatures extracted here. This is likely due to the solid angle coverage of the crystal, which is larger on the parts of the crystal closer to the source, which corresponds to lower photon energies. Therefore, the spectrometer has a higher collection efficiency at lower photon energies. This means that it appears as if there are more downshifted photons than there are, which skews the apparent detailed balance. At the same time, the $(E_{s}/E_{i})^{2}$ factor in Eq. (1) term mostly counteracts this effect by increasing the number of upshifted photons than detailed balance otherwise implies. While these effects will not be explicitly handled here since they mostly cancel, future improvements in analysis could account for them.
@@
Otherwise, the stability of the convergences and accuracy of the inferred temperature demonstrates the ITCF method can accurately extract system conditions. To within at least $\pm 250$ eV around the elastic feature, the convolution approximation appears to hold very well.
Figure 6: Examples of the ITCF for $T=20$ eV at scattering angle 1010 for different integration limits $x$. The location of the minima of each curve, which is used to infer temperature, are indicated by the circles. The minimum point of the exact temperature is indicated by the black dashed line.
@@
For the $T=15$ eV and 20 eV cases, sharp drops of the inferred temperature to $T_{\rm exact}/2$ are eventually seen as $x$ increases – such behaviour will also be seen later when using the ratios of ITCFs. These correspond to the ITCF losing its minimum point as the upshifted signal becomes very weak and few, if any, photons are measured. This causes the $\tau>\beta/2$ side of the ITCF to collapse, at which point the ITCF minimum just becomes the upper limit of the considered $\tau$-range, here restricted to $\tau_{\rm max}\ =\beta$. An example of this behaviour is shown for $T=20$ eV at scattering angle $10$ in Fig. 6, which plots the ITCFs for different integration limits. Such drops are not seen for $T=50$ eV and 100 eV since there is strong upshifted signal up to the edge of the detector.
These drops in the inferred temperature indicate the limits of being able to perform the deconvolution and extract temperature due to lack of upshifted signal. It also demonstrates that the lower temperature limit that the ITCF method can measure is ultimately determined by the photon statistics: enough signal needs to be measured sufficiently far from the elastic feature in both the upshifted and downshifted sides of the spectrum so that convergence in $x$ can be observed. The more extended the SIF is, the further away this point will be due to the need to deconvolve a broader function. And at the same time, the detailed balance relationship means the number of upshifted photons is exponentially reduced as the temperature decreases.
@@
Likewise, the ratio of inelastic to elastic scattering matters since it is the inelastic signal which needs to be measured in order to probe detailed balance. To emphasise this point, Fig. 5 shows the convergence of the inferred temperature for $T=15$ eV and 20 eV when the Rayleigh weight has been reduced by a factor of 4. In these cases, it takes a larger integration range $x$ before the ITCF loses its minimum, and the agreement between the temperatures at the different angles is improved. Convergences for a given number of input photons therefore have a strong dependence on the Rayleigh weight, and the models in MCSS are known to often overestimate the elastic scattering contribution [6, 5]. In the remainder of the article, the standard DSF model is used for the simulations, but will be compared to results for when the Rayleigh weight is reduced as well to understand the role of the elastic feature in extracting the temperature.
@@ III.2 Temperature from ITCF Ratios
For the $T=20$ eV, 50 eV, and 100 eV cases, there are three ratios which show convergence, all with the higher scattering angle in the numerator of the ITCF ratio. The $T=50$ eV and 100 eV cases in particular produce very accurate temperatures over a very large range of $x$ due to their large upshifted signal. We also note these two high temperatures show more accurate temperature inferences using the ratio method in Fig. 7 than using the standard ITCF method in Fig. 4. The other ratios do not show meaningful, if any, convergence – however, these can simply be discarded since their inverses do show convergence (or, equivalently, one could look at their maximum points). For $T=20$ eV, at $x>160$ eV, there are drops in the inferred temperature to $T_{\rm exact}/2$ which, as with ITCFs for Fig. 4, indicates the ITCF ratio has lost its well-defined minimum.
@@
For $T=15$ eV, only one ITCF ratio, 201020/10, shows a region of convergence in its minimum that could be used to extract temperature; the remainder do not yield useful information. This is because of the very low number of upshifted photons measured. Indeed, the situation is worse than in Fig. 4 since we now require upshifted signal in two ITCFs in order to measure a temperature. If the Rayleigh weight is reduced by a factor of 4, the ratio method is more successful at extracting the temperature, as shown in Fig. 8. In this case, there are three ratios – 2015020/150, 15010150/10, and 201020/10 – which have a region of convergence in their minima that can infer the temperature. When $x>140$ eV, these minima of these ratios begin to diverge or collapse to $T_{\rm exact}/2$. The 15 eV results again highlight that the ability to measure low temperatures is strongly determined by the number of measured upshifted photons, which from detailed balance becomes exponentially more difficult as the temperature decreases. For the HAPG crystal used here, the problem is made worse by the extended crystal IF, meaning upshifted photons need to be measured relatively far from the elastic peak in order to observe convergence. Low-mosaicity HOPG crystals may therefore be more useful for low temperature measurements since the Gaussian distribution of crystallites reduces the extent of the crystal broadening [36].
@@
Figs. 7 and 8 also highlight the strength of collecting XRTS spectra at three or more scattering angles: the analysis of the convergence is made much simpler by being able to cross-check the extracted temperatures between three or more ratios. At the same time, for the T=20T=20–100 eV cases, the high consistency in the inferred temperatures suggests that discrepancies of even a few eV in the inferred temperatures from the ratios of three or more scattering angles would be a strong indication of non-equilibrium effects in the system [77, 2]. However, as indicated by the $T=15$ eV results, the ability to do so is closely tied to the number of measured photons, and this caveat is explored in more detail in the following Section III.3.